# EZIGen: Enhancing zero-shot subject-driven image generation with precise subject encoding and decoupled guidance

### v1

@@ 3.3 Iterative Appearance Transfer
Based on the above analysis and empirical results, we observe that the initial layout can influence the final subject-driven generation outcomes. If the subject in the initial layout is similar to the subject image, the transfer process improves, as the adaptor can more easily establish correspondence within the noisy latent space produced by the Main UNet.
Therefore, we introduce an iterative generation scheme. Specifically, after each transfer process, the generation is restarted by treating the generated image as a new noisy layout. Noise is then added to revert the clear image back to $x^{\prime}$ according to timestep $T_{\text{layout}}$, allowing further editing while preserving the layout, as shown in Fig. 5, line 4. This process continues iteratively until a satisfactory result is achieved.
With the decoupling operation and iterative generation process, the model clearly delineates the role of each guidance in manipulating the generation process and also ensures comprehensive subject identity injection. This approach achieves a better balance between subject identity preservation and text prompt coherence without compromising either aspect.
@@ 3.4 Subject-Driven Image Editing
We discover that the Appearance Transfer Process can naturally function as an effective subject-driven image editor. This is achieved by integrating an object mask and replacing noise addition with image inversion (Lu et al. (2023)), which converts the generated image back into the layout latent, preserving the background. Specifically, similar to the noise addition described in Sec. 3.3, given a real image, we first partially invert it based on timestep $T_{\text{layout}}$ to obtain a coarse layout latent $x^{\prime}\text{r}$. Next, we initiate the iterative appearance transfer process to inject the subject feature, resulting in an incomplete edited latent $\hat{x}^{\prime}\text{r}$. To maintain the original background, we separate the edited foreground from $\hat{x}^{\prime}\text{r}$ and the background from $x^{\prime}\text{r}$ using a user-provided foreground mask $M$:
@@ 4 Experiment
4.1 Evaluation and Implementation Details
Benchmark. We conduct both qualitative and quantitative evaluation on three public benchmarks, namely DreamBench (Ruiz et al. (2023)) for subject-driven image generation, DreamEdit (Li et al. (2023)) for subject-driven image editing, and FastComposerBench (Xiao et al. (2023)) for human content generation. For evaluation on DreamBench and DreamEdit, we follow Subject Diffusion (Ma et al. (2023)) to calculate the average score among 6 random runs. For human content generation, We follow FastComposer to use its official API for evaluation.
Evaluation Metrics. For the experiment on DreamBench, we follow previous methods to report CLIP-T, CLIP-I score, and DINO score to evaluate text adherence and subject identity preservation, we follow (Ma et al. (2023)) to take CLIP-T and DINO scores as main evaluation metrics. Regarding subject-driven editing, we follow (Li et al. (2023)) to evaluate DINO/CLIP-sub score for foreground object similarities and DINO/CLIP-back score for background respectively, we leverage Grounded-SAM to obtain subject fore-/back-ground mask during evaluation for score calculation. For the human content generation task, we follow FastComposer (Xiao et al. (2023)) to calculate ID preservation and Prompt Consistency.
Training dataset construction. We build training image pairs from the COCO2014 and YoutubeVIS datasets. In the COCO2014 dataset, for each target image, we randomly crop out 1 to 4 objects as the subject images according to the segmentation masks; whereas in YoutubeVIS, we follow (Chen et al. (2024)) to extract the image of the same subject from different frames as subject-target pairs. We sample 100k image pairs for each dataset, adding up to 200k image pairs in total.
Implementation details. We leverage Stable Diffusion V2.1-base as the base model. All input and output image resolutions are fixed to 512$\times$512. During training, we initialize the adapter from the self-attention module within each transformer block, and we train the model for 1 epoch with the training batch size set to 1, using Adam optimizer and learning rate $51e-$ 5. During inference, the iteration will automatically stop when the generated image shares a high similarity with the image from the previous loop. Besides, we further limits a broader range for the iteration steps with a minimum step of 3 and a maximum of 10.
@@ 4.1 Evaluation and Implementation Details
Figure 6: Qualitative comparison with existing subject-driven generation methods. Since Subject Diffusion is not publically available, its resultant images are directly taken from the original paper.
@@
Table 3: Scores on DreamEditBench. Our design outperforms previous methods w.r.t. foreground subject ID preservation, highlighting the effectiveness of our subject feature encoding and injecting strategy. Results with $\ddagger$ are referenced from DreamEdit.
@@
Table 4: Performances on single-subject human image generation without training on any large-scale or domain-specific dataset compared to previous methods. Results with $\ddagger$ are referenced from Subject Diffusion.
Figure 7: Qualitative results on DreamEditBench with existing subject-driven editing methods1.
Figure 8: Qualitative evaluation on single-subject human image generation.
@@ 4.2 Comparing with existing subject-driven generation methods
In Tab. 2, we compare our method with existing zero-shot subject-driven approaches on the DreamBench dataset (Ruiz et al. (2023)), including tuning-based methods for reference. We also present the number of reference images required for each subject and the training dataset size to highlight the cost of each method. The overall results show that our base model achieves state-of-the-art performance text-following, subject identity preservation, and balance among all methods. Compared to the previous state-of-the-art, Subject Diffusion (Ma et al. (2023)), our model outperforms in both CLIP-T (0.315 vs. 0.293) and DINO score (0.718 vs. 0.711), demonstrating superior flexibility in generated images and higher subject identity preservation, without sacrificing balance. Moreover, this was accomplished with 100$\times$ less training data, which we attribute to our simplified design and the use of the Reference UNet. These results highlight the advantages of using a Reference UNet over CLIP for feature extraction, alongside our decoupling design that improves text coherence. When compared with BootPIG (Purushwalkam et al. (2024)), our method surpasses it in text-following ability (0.316 vs. 0.311) and significantly outperforms in DINO score (0.716 vs. 0.628), while requiring fewer subject images. This is due to our more advanced utilization of the Reference UNet. Additionally, our method outperforms previous open-source zero-shot methods (Wei et al. (2023); Li et al. (2024a)) across all evaluation metrics. 00footnotetext: 1 We follow DreamEdit to take the result images from ImageHub evaluation platform.
4.3 Valitation on subject-driven editing task.
As outlined in Sec. 3.4, our method is well-suited for subject-driven image editing by replacing noise addition in the Appearance Transfer Process with image inversion. We compare our model with previous state-of-the-art methods on the DreamEdit benchmark, as shown in Tab. 4 and Fig. 7. Using DINO-sub as the primary metric for subject fidelity, our method significantly outperforms previous approaches. We attribute this improvement to our advanced subject feature extraction paradigm and image Adapter design, where the Reference UNet extracts high-quality features that preserve exceptional subject details and the Adapter matches the appearances accurately. In addition, our method achieves superior background preservation (DINO-sub and CLIP-sub), which we attribute to the explicit separation of the foreground and background using a foreground mask. This approach ensures that inversion errors and pixel alterations are confined to the foreground, while the background remains unaffected throughout the editing process.
@@
We also demonstrate the effectiveness of our method for subject-driven human image generation. Unlike FastComposer (Xiao et al. (2023)) and InstantBooth (Shi et al. (2024)), which were trained on domain-specific datasets, or Subject Diffusion (Ma et al. (2023)), which involves large-scale pretraining, our model achieves high-quality human content generation using only a normal-scale open-domain dataset. As shown in Tab. 4, our approach delivers competitive results based on the evaluation protocols introduced by FastComposer. Specifically, in terms of subject ID preservation, our performance is on par with Subject Diffusion and outperforms FastComposer. However, for prompt consistency, while our method lags behind FastComposer and DreamBooth, it still surpasses Subject Diffusion. These results can also be reflected in Fig. 8, where we qualitatively compare our results with those of Subject Diffusion and FastComposer. Our method captures accurate facial details, such as shape, hair, and facial landmarks. FastComposer excels in generating more versatile and stylized content, which contributes to its higher prompt consistency scores.
@@ 4.5 Ablation Study
We conducted ablation studies to evaluate the effects of various components in our method, including the Subject Encoding and Injection module, the naive single-stage Appearance Transfer, and the Iterative Appearance Transfer process. The quantitative and qualitative results are reported in Tab. 5 and Fig. 5.
Effectiveness of UNet-based Subject Encoding and Injection. To fully evaluate the Adapter’s appearance transfer ability, we use a totally noised layout image (or random noise) as a starting point, allowing the Adapter to transfer subject identities at all timesteps. As shown in Tab. 5 line 1 and Fig. 5 line 2, the Adapter demonstrates strong identity preservation, achieving a DINO score of 0.762 and a CLIP-I score of 0.808. However, this brute-force injection overrides text guidance in the early timesteps, resulting in a lower CLIP-T score of 0.286.
  Encode & Inject Decoupled Generation Process Iterative Appearance Transfer CLIP-T CLIP-I DINO
0.286 0.808 0.762
0.302 0.743 0.663
0.316 0.782 0.718
@@
Decoupling the Layout Generation Process. This experiment confirms the effectiveness of separating the Layout Generation Process from Appearance Transfer. As seen in Tab. 5 line 2 and Fig. 5 line 3, the CLIP-T score improves significantly (from 0.286 to 0.302), as this design better maintains the layout created by the text prompt. However, when there is a large discrepancy between the layout and the subject, the appearance transfer can be incomplete, leading to artifacts and blurs, as seen in the final image of Fig. 5 line 3. This results in reduced subject similarity scores (DINO/CLIP-I).
Effectiveness of Iterative Appearance Transfer. The iterative approach successfully transfers the subject’s appearance (a small brown dog) into the generated image after several iterations, as shown in Fig. 5 line 4. This leads to substantial improvements in all scores: CLIP-T increases from 0.302 to 0.316, CLIP-I from 0.743 to 0.782, and DINO score from 0.663 to 0.718. This demonstrates that the iterative process resolves the artifacts from the Layout Generation Process and also produces a complete and accurate image.
@@ 4.6 Superiority of our Reference UNet design among other alternatives
This section evaluates the effectiveness and efficiency of our Reference UNet subject extractor in comparison to CLIP, DINO, and alternative Reference UNet configurations. We use three baseline methods—Subject Diffusion (Ma et al. (2023)), AnyDoor (Chen et al. (2024)), and BootPIG (Purushwalkam et al. (2024))—as the current best practices for feature extraction. To ensure a fair comparison, we ablate the models for AnyDoor and Subject Diffusion (lines 2 and 4) so that subject information is derived solely from the feature extractor. The results are presented in Tab. 7
Comparing with CLIP- and DINO-based methods. First, to evaluate the efficiency of using Reference UNet versus CLIP and DINO, we directly replaced our “Reference UNet + Adapter” combination with “CLIP/DINO + Projector + Adapter” and compared performances under identical training settings. As shown in Tab. 7 lines 1 and 3, this replacement led to trivial solutions, as the Adapter struggled to establish attention between misaligned feature maps, causing the model to rely primarily on the text prompt for reconstruction. To address this, additional regularization is typically required to overcome feature space misalignment and force the model to follow the subject guidance. For example, Subject Diffusion employs location control as regularization, while AnyDoor replaces text tokens with image tokens and crops a scene image for the subject feature to fill. However, even with regularization, the required dataset sizes remain substantial, at approximately 76M and 9.6M image pairs, respectively. In contrast, our method completes training with only 0.2M image pairs.
Second, to validate the effectiveness of our design over CLIP and DINO extractors in encoding subject information, we report the DINO and CLIP-I scores on the DreamBench dataset using the same evaluation protocol. As shown in Tab. 7 lines 2 and 4, and Tab. 7, our Reference UNet significantly outperforms both DINO- and CLIP-based configurations. This suggests that, although CLIP and DINO are highly effective for high-fidelity image representation across various downstream tasks (Liu et al. (2024); Zhou et al. (2023); Sun et al. (2024b)), the UNet architecture is inherently more suitable for providing latent representations for high-quality generation, as it was specifically designed for this purpose.
@@
DINO 1) DINO + Projector 0.661 0.421 0.2M
2) AnyDoor w/o high-freq map $\dagger$ 0.775 0.710 $\sim$9M
CLIP 3) CLIP + Projector 0.682 0.433 0.2M
4) Subject Diff. w/o image CLS token $\dagger$ 0.719 0.637 $\sim$76M
Fixed Ref UNet 5) Ours w/ all-step transfer 0.808 0.762 0.2M
Table 6: Comparing subject extractor designs. Results with $\dagger$ are taken from the original paper.
Table 7: Comparison with state-of-the-art CLIP-based and DINO-based methods Subject Diffusion (Ma et al. (2023)) and AnyDoor (Chen et al. (2024)). Our method maintains high-fidelity subject appearance and a more flexible pose variation.
  Train Ref. UNet Ref. UNet Timestep CLIP-I DINO
1) Adaptive 0.764 0.589
2) Adaptive 0.771 0.615
3) $T_{\text{sub}}$=1 0.778 0.668
4) $T_{\text{sub}}$=800 0.427 0.427
5) $T_{\text{sub}}$=500 0.679 0.598
6) $T_{\text{sub}}$=300 0.789 0.661
7) $T_{\text{sub}}$=100 0.789 0.725
8) Ours $T_{\text{sub}}$=1 0.808 0.762
Figure 9: Evalutating subject ID preservation under different UNet training strategies. As shown in the table on the left, we obtain the best result when we fix the Reference UNet parameters and assign it with a small denoising timestep $T_{\text{sub}}$. On the right are the qualitative comparisons accordingly.
Evaluating alternative Reference UNet configurations. We also examined the impact of tuning Reference UNet parameters and adding varying levels of noise to subject images on simulating the denoising process, as shown in Fig. 9. The table on the left lists the experiment settings and their corresponding performances, where check marks/cross marks indicate whether the Reference UNet is trained or frozen. “Adaptive” means that the Reference UNet shares the same denoising timestep as the Main UNet, while arguments starting with “$T_{\text{sub}}$” refer to the addition of a fixed noise according to timestep $T_{\text{sub}}$ to the subject image.
From lines 1 and 2, we observe that tuning the Reference UNet can disrupt the original denoising capabilities of the Stable Diffusion UNet, leading to a drop in performance compared to freezing the Reference UNet. Comparing lines 1 and 3, we find that fixing the denoising timestep of the Reference UNet to a small $T_{\text{sub}}$ yields better subject identity with more details and clarity, as this approach ensures clear feature maps. Further, we validated the importance of a small $T_{\text{sub}}$ by fixing the Reference UNet and adjusting $T_{\text{sub}}$. As demonstrated in lines 4 to 8, subject identity scores steadily increase as $T_{\text{sub}}$ decreases. These results follow the intuition that the Reference UNet tends to encode clearer features in the last timesteps, as is in the generation process.

### v2

@@ 3.3 Iterative Appearance Transfer
Based on the above analysis and empirical results, we observe that the initial layout can influence the final subject-driven generation outcomes. When the initial layout resembles the subject image, the transfer process improves as the adapter establishes better correspondence within the Main UNet’s noisy latent space. To enhance this, we introduce an iterative generation scheme: after each transfer, the generated image becomes the new noisy layout, with noise added according to timestep $T_{\text{layout}}$ for further editing, as shown in the bottom part of Fig. 4. This process repeats until the similarity between the newly generated image and the previous image exceeds a predefined threshold, indicating that the appearance transfer is complete and no further information is added.
Integrating the designs above, our model successfully balances subject identity preservation and text adherance, ensuring comprehensive guidance-following without compromise.
@@ 3.4 Subject-Driven Image Editing
We discover that the Appearance Transfer Process can naturally function as an effective subject-driven image editor. This is achieved by integrating an object mask and replacing noise addition with image inversion (Lu et al. (2023)), which converts the generated image back into the latent layout, preserving the background. Specifically, similar to the noise addition described in Sec. 3.3, given a real image, we first partially invert it based on timestep $T_{\text{layout}}$ to obtain a coarse layout latent $x^{\prime}_{\text{r}}$. Next, we initiate the iterative appearance transfer process to inject the subject feature, resulting in an incomplete edited latent $\hat{x}^{\prime}_{\text{r}}$. To maintain the original background, we separate the edited foreground from $\hat{x}^{\prime}_{\text{r}}$ and the background from $x^{\prime}_{\text{r}}$ using a user-provided foreground mask $M$:
@@ 4 Experiment
4.1 Implementation Details
Benchmark and Evaluation. We evaluate our design on three benchmarks: DreamBench (Ruiz et al. (2023)) for subject-driven image generation, DreamEdit (Li et al. (2023)) for subject-driven image editing, and FastComposerBench (Xiao et al. (2023)) for human content generation. For DreamBench and DreamEdit, we follow Subject Diffusion’s protocol, averaging scores over 6 random runs. For human content generation, we use the FastComposer API. For DreamBench, we report CLIP-T, CLIP-I, and DINO scores to assess text adherence and subject identity. For subject-driven editing, we evaluate DINO/CLIP-sub for foreground similarities and DINO/CLIP-back for background, using SAM (Kirillov et al. (2023)) for mask extraction. For the aforementioned tasks, we take the DINO-related scores as the main metric for subject similarity, as it better depicts the detailed patch-level differences between images. Finally, for human content generation, we calculate ID preservation and prompt consistency, following Xiao et al. (2023).
Training Dataset Construction. We create training pairs from COCO2014 and YoutubeVIS. In COCO2014, 1-4 objects are cropped from target image as subject images, while in YoutubeVIS, we extract images of the same subject from different frames as pairs, following Chen et al. (2024). We sampled 100k pairs from each dataset, totaling 200k pairs.
Experiment settings. We utilize Stable Diffusion V2.1-base, with the image resolution fixed at 512×512 for all experiments. The adapter is initialized from the self-attention module within each transformer block to enhance compatibility. We train the model for 1 epoch using a batch size of 1, with the Adam optimizer and a learning rate of $51e-$ 5. During inference, iterations are designed to stop automatically when the newly generated image exhibits a sufficiently high similarity with the image from the previous loop, ensuring efficient convergence and maintaining generation quality.
@@ 4.1 Implementation Details
Figure 5: Comparison with existing subject-driven generation methods. Since Subject Diffusion is not publically available, we take resultant images from the original paper.
@@
Table 2: Scores on DreamEditBench. The scores show the effectiveness of our subject encoding. Results with $\ddagger$ are referenced from DreamEdit.
@@
Table 3: Performances on single-subject human image generation. Results with $\ddagger$ are referenced from Subject Diffusion.
Figure 6: Comparing on 1) subject-driven editing1 and 2) human content generation tasks.
@@ 4.2 Comparing with existing subject-driven generation methods
In Tab. 1, we compare our method with existing zero-shot subject-driven approaches on the DreamBench dataset (Ruiz et al. (2023)), including tuning-based methods for reference. We also present the number of reference images required for each subject and the training dataset size to highlight the cost of each method. The overall results show that our base model achieves state-of-the-art performance text-following, subject identity preservation, and balance among all methods. Compared to the previous state-of-the-art, Subject Diffusion (Ma et al. (2023)), our model outperforms in both CLIP-T (0.316 vs. 0.293) and DINO score (0.718 vs. 0.711), demonstrating superior flexibility in generated images and higher subject identity preservation, without sacrificing balance. Moreover, this was accomplished with 100$\times$ less training data, which we attribute to our simplified design and the use of the Reference UNet. These results highlight the advantages of using a Reference UNet over CLIP for feature extraction, alongside our decoupling design that improves text coherence. When compared with BootPIG (Purushwalkam et al. (2024)), our method surpasses it in text-following ability (0.316 vs. 0.311) and significantly outperforms in DINO score (0.718 vs. 0.674), while requiring fewer subject images. This is due to our more advanced utilization of the Reference UNet. Additionally, our method outperforms previous open-source zero-shot methods (Wei et al. (2023); Li et al. (2024a)) across all evaluation metrics. 00footnotetext: 1 We follow DreamEdit to take the result images from ImageHub evaluation platform.
4.3 Validation on Subject-Driven Editing Task
As outlined in Sec. 3.4, our method adapts well to subject-driven image editing by replacing noise addition with image inversion (Lu et al. (2023)) in the Appearance Transfer Process. We compare our model against previous state-of-the-art methods on the DreamEdit benchmark, as shown in Tab. 3 and Fig. 6(1). Using DINO-sub as the main metric for subject fidelity, our method significantly outperforms others, primarily due to our advanced subject feature utilization, where the Reference UNet extracts high-quality, detailed features, and the Adapter accurately matches subject appearances to the layout. Additionally, our approach achieves superior background preservation, thanks to the explicit separation of foreground and background, which effectively confines inversion errors and pixel alterations strictly to the foreground, leaving the background entirely intact.
@@ 4.3 Validation on Subject-Driven Editing Task
We demonstrate the effectiveness of our method for subject-driven human image generation. Unlike FastComposer (Xiao et al. (2023)), which relies on domain-specific datasets, or Subject Diffusion (Ma et al. (2023)), which requires large-scale pretraining, our model achieves high-quality results using only a normal-scale open-domain dataset. As shown in Tab. 3, our approach delivers very competitive results. In terms of subject ID preservation, we match Subject Diffusion and outperform FastComposer. While our prompt consistency lags behind FastComposer and DreamBooth, it still exceeds Subject Diffusion. These findings are also reflected in Fig. 6(2), where our method accurately captures facial details, such as shape, hair, and facial landmarks.
@@ 4.5 Ablation Study
We conducted ablation studies to evaluate the effects of various components in our method, including the Subject Encoding and Injection module, the naive single-stage Appearance Transfer, and the Iterative Appearance Transfer process. The results are reported correspondingly in Tab. 4 and Fig. 7, with markings from 1) to 4).
Effectiveness of UNet-based Subject Encoding and Injection. To evaluate the quality of the encoded subject feature, we start with a fully noised layout image (or random noise) and apply subject guidance at all timesteps. As shown in experiments 1) and 2), compared to the original text-guided generation, the encoded subject representation demonstrates strong identity preservation, achieving a DINO score of 0.762 and a CLIP-I score of 0.808. However, the strong subject features tend to override the weaker layout latent in early timesteps, leading to serve copy-paste effect and a lower CLIP-T score of 0.286.
  #ID Encode & Inject Decoupled Generation Process Iterative App. Transfer CLIP-T CLIP-I DINO
1 0.327 0.658 0.362
2 0.286 0.808 0.762
3 0.321 0.714 0.560
4 0.316 0.782 0.718
@@
Figure 7: Qualitative result of the ablation study.
Decoupling the generation process. The experiment in 3) confirms the effectiveness of separating the Layout Generation Process from Appearance Transfer, in which the CLIP-T score improves significantly from 0.286 to 0.321, as this design explicitly maintains the layout. However, when there is a large discrepancy between the layout content and the subject, the appearance transfer can be incomplete, resulting in reduced subject similarity scores compared to experiment 1), where the DINO score drops from 0.762 to 0.560.
Effectiveness of Iterative Appearance Transfer. Experiment 4) indicates a successful appearance transfer after introducing the iterative process. Such gentle refinement to the incomplete generated image leads to substantial improvements in subject appearance similarity: CLIP-I improved from 0.714 to 0.782, DINO score from 0.560 to 0.718, while the text-guided semantics of the generation images are slightly disrupted during the iterative process, and thus the CLIP-T score drop marginally from 0.321 to 0.316, our model reaches the best balance between two guidance under this setting.
@@ 4.6 Superiority of our Reference UNet design among other alternatives
We evaluate the effectiveness and efficiency of our Reference UNet subject extractor compared to CLIP, DINO, and alternative Reference UNet configurations. We use three baseline methods—Subject Diffusion (Ma et al. (2023)), AnyDoor (Chen et al. (2024)), and BootPIG (Purushwalkam et al. (2024))—as the current best practices for feature extraction. To ensure a fair comparison, we ablate the models for AnyDoor and Subject Diffusion (lines 2 and 4) so that subject information is derived solely from the feature extractor.
Comparing with CLIP- and DINO-based methods. First, to evaluate the efficiency of using Reference UNet versus CLIP and DINO, we directly replaced our “Reference UNet + Adapter” combination with “CLIP/DINO + Projector + Adapter” and compared performances under identical training settings. As shown in Tab. 5 lines (a) and (c), this replacement led to trivial solutions, as the Adapter struggled to establish attention between misaligned feature maps, causing the model to rely primarily on the text prompt for reconstruction. To address this, additional regularization is typically required to overcome feature space misalignment and force the model to follow the subject guidance. For example, Subject Diffusion employs location control as regularization, while AnyDoor replaces text tokens with image tokens and crops a scene image for the subject feature to fill. However, even with regularization, the required dataset sizes remain substantial, at approximately 76M and 9M image pairs, respectively. In contrast, with closer feature spaces, our method completes training with only 0.2M image pairs. Second, to validate the effectiveness of our design over CLIP and DINO extractors in encoding subject information, we report the DINO and CLIP-I scores on the DreamBench dataset using the same evaluation protocol, as shown in Tab. 5 lines (b) and (d), and Fig. 8, our Reference UNet significantly outperforms both DINO- and CLIP-based configurations. This suggests that, although CLIP and DINO are highly effective for high-fidelity image representation across various downstream tasks (Liu et al. (2024); Zhou et al. (2023); Sun et al. (2024b)), the Stable Diffusion UNet is more suitable for providing features for high-quality image generation, as it was specifically designed.
@@
DINO a) DINO + Projector 0.661 0.421 0.2M
b) AnyDoor w/o high-freq map $\dagger$ 0.775 0.710 $\sim$9M
CLIP c) CLIP + Projector 0.682 0.433 0.2M
d) Subject Diff. w/o image CLS token $\dagger$ 0.719 0.637 $\sim$76M
Fixed Ref UNet e) Ours w/ all-step transfer 0.808 0.762 0.2M
Table 5: Comparing extractor designs. Results with $\dagger$ are taken from the original paper.
Figure 8: Comparison with state-of-the-art CLIP-based and DINO-based methods Subject Diffusion (Ma et al. (2023)) and AnyDoor (Chen et al. (2024)).
  #ID Train Ref. UNet Ref. UNet Timestep CLIP-I DINO
a) BootPIG* Adaptive 0.764 0.589
b) Adaptive 0.771 0.615
c) $T_{\text{sub}}$=1 0.778 0.668
d) $T_{\text{sub}}$=800 0.427 0.427
e) $T_{\text{sub}}$=500 0.679 0.598
f) $T_{\text{sub}}$=300 0.789 0.661
g) $T_{\text{sub}}$=100 0.789 0.725
h) Ours $T_{\text{sub}}$=1 0.808 0.762
Figure 9: Evalutating subject ID preservation under different UNet configurations. We obtain the best result when we fix the Reference UNet parameters and assign it with a small denoising timestep $T_{\text{sub}}$. “BootPIG*” indicates BootPIG’s training strategy.
Evaluating alternative Reference UNet configurations. We examined the impact of tuning Reference UNet parameters and adding varying levels of noise to subject images, as shown in Fig. 9. The table on the left lists the experiment settings and their performances, where check marks/cross marks indicate whether the Reference UNet is trained or frozen. “Adaptive” means that the Reference UNet shares the same denoising timestep as the Main UNet, while arguments starting with ”$T_{\text{sub}}$” indicate the level of noise we add to the subject image. From lines (a) and (b), we find that tuning the Reference UNet disrupts the original denoising capabilities of Stable Diffusion, resulting in a performance drop compared to freezing it. Lines (a) and (c) show that fixing the denoising timestep of the Reference UNet to a small $T_{\text{sub}}$ produces better subject identity with clearer feature maps. By fixing the Reference UNet and adjusting $T_{\text{sub}}$, we validated its importance; as shown in lines (d) to (h), subject identity scores increase as $T_{\text{sub}}$ decreases. This aligns with the intuition that the Reference UNet encodes clearer features in the final timesteps of the generation process.